a grocer plans to mix candy that sells for $1.20 a pound with candy that sells for $2.40 a pound to get a mixture that he plans to sell for $1.65 a pound. how much of the $1.20 and the $2.40 candy should he use if he wants 40 pounds of the mix?
Answer by jorel1380(3719)

1.20p+2.40(40-p)=40(1.65)
1.20p+96-2.40p=66
1.20p=30
p=30/1.20=25
The grocer needs 25 lbs. of $1.20 candy and 15 lbs. of $2.40 candy..